OpenStudy (anonymous):

8 is to 64 as 2 is to x

OpenStudy (anonymous):

How do I solve to find x?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Can you explain how you got 4?

OpenStudy (anonymous):

well you multiply 8 how many times to get to 64 then multiply two by how many times it took 8 to get to 64 so 8 times 8 is 64 so you multiply 2 times 8 to get 16 so there is your answer

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you take squares

OpenStudy (anonymous):

8^2= 64 2^2=4

OpenStudy (anonymous):

well either would work

OpenStudy (anonymous):

but you dont get 16

OpenStudy (anonymous):

right now im lost guys

OpenStudy (anonymous):

Okay to clarify

OpenStudy (anonymous):

8^2=64 and 2^2=4, the pattern is squaring, which means multiplying numbers by the same number

OpenStudy (anonymous):

8*8=64 2*2=4

OpenStudy (anonymous):

trust me i did this 2 years ago

OpenStudy (anonymous):

so the x would = 4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you again

OpenStudy (anonymous):

well ok then